# Hamiltonian reductions as affine closures of cotangent bundles

### v1

@@ Abstract.
For an irreducible non-singular affine $G$-variety $Y$ whose action is $2$-large, we prove that the Hamiltonian reduction $T^{*}Y/\!\!/\!\!/G$ is a symplectic variety with terminal singularities, isomorphic to the affine closure of $T^{*}Z_{\operatorname{reg}}$ for $Z\coloneqq Y/\!/G$. As applications, we provide a short proof of G. Schwarz’s theorem on the graded surjectivity of the push‑forward map $\mathcal{D}(Y)^{G}\rightarrow\mathcal{D}(Z)$, and we establish the surjectivity of the symbol map on $Z$.
@@ 1. Introduction
Recently, we started a systematic investigation of a geometric construction of symplectic varieties through cotangent bundles. Namely for a non-singular variety $Y$ with $T^{*}Y$ its cotangent bundle, denote by $\mathcal{O}(T^{*}Y)$ the ring of regular functions on $T^{*}Y$, then we consider the affinization morphism
@@
When $Y$ is a non-singular projective variety, it is shown in [FL25b] that $\mathcal{Z}_{Y}$ is a symplectic variety if the tangent bundle $T_{Y}$ is big and if the projective bundle $\mathbb{P}T_{Y}$ is of Fano type. The most interesting case is where $Y$ is quasi-affine but not affine. In this case, $\psi_{Y}$ is just an open embedding and, therefore, $\mathcal{Z}_{Y}$ is the affine closure of $T^{*}Y$. It remains a challenging problem to determine when $\mathcal{Z}_{Y}$ is a symplectic variety.
Another geometric way to construct symplectic varieties is through Hamiltonian reductions, which is a powerful tool in both physics and geometry to produce new symplectic manifolds from a Hamiltonian action on a symplectic manifold. To be more precise, let $(X,\omega)$ be a non-singular affine symplectic variety on which there exists a Hamiltonian action by a reductive group $G$ so that the associated moment map $\mu:X\to\mathfrak{g}^{*}$ is $G$-equivariant (see § 2). We denote by $N_{X}$ the scheme-theoretic fiber $\mu^{-1}(0)$, called the shell. The Hamiltonian reduction X///GX/\!\!/\!\!/G is by definition the affine GIT quotient:
@@
Let $G$ be a reductive group and $Y$ an irreducible non-singular affine $G$-variety which is $2$-large. Then the following hold:
@@
    The Hamiltonian reduction $T^{*}Y/\!\!/\!\!/G$ is the affine closure of $T^{*}(Y/\!/G)_{\rm reg}$.
The $2$-largeness condition in Theorem 1.1 is not overly restrictive: as shown in [Sch95, Corollary 11.6 (2)] (see also [HSS20, Theorem 3.6]), most $G$-modules are in fact $2$-large, with only finitely many exceptions under mild assumptions.
We next investigate the ($\mathbb{Q}$-)factoriality of the Hamiltonian reduction $T^{*}Y/\!\!/\!\!/G$ and use this to deduce the non‑existence of symplectic resolutions under certain conditions. Recall that the abelianization of $G$ is $G_{\operatorname{ab}}\coloneqq G/[G,G]$. The group $G$ is called perfect if $G_{\operatorname{ab}}=\{e\}$, i.e., $G=[G,G]$.
Theorem 1.2.
Let $G$ be a reductive group and $Y$ an irreducible non-singular affine $G$-variety which is $2$-large. Assume that either $\mathcal{O}(Y)^{\times}=\mathbb{C}^{\times}$ or $G$ is connected.
@@
    If $G_{\operatorname{ab}}$ is finite and $\operatorname{Pic}(Y)$ is finite, then $\operatorname{Cl}(T^{*}Y/\!\!/\!\!/G)$ is finite. In particular, $T^{*}Y/\!\!/\!\!/G$ is $\mathbb{Q}$-factorial, and if $T^{*}Y/\!\!/\!\!/G$ is singular, it admits no symplectic resolution.
  2. (1.2.2)
For the special case in which $Y$ is a linear representation of $G$ that is $3$-large (a condition stronger than $2$-largeness), part (1.1.1) of Theorem 1.1 was proved in [HSS20] (see also [HSS24a, HSS24b] for other related results), while the local factoriality and $\mathbb{Q}$-factoriality of $T^{*}Y/\!\!/\!\!/G$ were established in [BS19, Theorem 1.2]. Furthermore, we present a partial converse of Theorem 1.2 in Corollary 4.7, and provide examples illustrating the sharpness of the above theorems in § 4.3.
@@
Let $G$ be a reductive group and $Y$ an irreducible non-singular affine $G$-variety which is $2$-large. Set $Z\coloneqq Y/\!/G$ with quotient map $\pi_{Y}\colon Y\rightarrow Z$.
@@ 2. Quasi-Hamiltonian varieties
Throughout this paper, we work over the complex numbers, and $G$ denotes always a complex reductive algebraic group. By varieties we mean separated reduced schemes of finite type over $\mathbb{C}$, which are not necessarily irreducible.
@@ 2.1. Hamiltonian actions
Then $X$ is called a Hamiltonian $G$-variety so that $V_{\xi}$ and $H_{\xi}$ are called the Hamiltonian vector field and the Hamiltonian function associated to $\xi$, respectively.
For a Hamiltonian action $G$ on $X$, we can define a morphism $\mu_{G,X}\colon X\rightarrow\mathfrak{g}^{*}$ by the following formula:
@@
Given a quasi-Hamiltonian $G$-variety, since $X$ is normal and $\operatorname{codim}(X\setminus X_{\operatorname{reg}})\geq 2$, the moment map $\mu_{G,X_{\operatorname{reg}}}\colon X_{\operatorname{reg}}\rightarrow\mathfrak{g}^{*}$ extends uniquely to a $G$-equivariant morphism $X\rightarrow\mathfrak{g}^{*}$, denoted again by $\mu_{G,X}$. We shall call $\mu_{G,X}$ also the moment map of the quasi-Hamiltonian $G$-variety $X$.
@@
Let $Y$ be an irreducible non-singular $G$-variety, $X\coloneqq T^{*}Y$ the cotangent bundle of $Y$. Then $X$ is a non-singular symplectic variety with the canonical symplectic form $d\eta$, where $\eta$ is the Liouville $1$-form on $T^{*}Y$. Moreover, the lifted action of $G$ on $X$ is Hamiltonian and the moment map is given by
@@
where $y\in Y$, $\alpha\in T_{y}^{*}Y$ and $\xi\in\mathfrak{g}$. More generally, one can also construct singular affine Hamiltonian $G$-varieties by considering the affine closure of cotangent bundles. We refer the reader to [FL25a, FL25b] for examples.
@@ 2.2. The shell
Let $X$ be a quasi-Hamiltonian $G$-variety. The shell $N_{X}$ is defined to be the scheme-theoretic fiber of the moment map $\mu\colon X\rightarrow\mathfrak{g}^{*}$ over $0\in\mathfrak{g}^{*}$. We denote by $N_{X,\operatorname{fi}}$ the subset of $N_{X}$ consisting of points $x\in N_{X}\cap X_{\operatorname{reg}}$ such that the isotropy subgroup $G_{x}$ is finite. If $X$ is clear from the context, we will simply denote $N_{X}$ by $N$ and $N_{X,\operatorname{fi}}$ by $N_{\operatorname{fi}}$.
Lemma 2.6 (cf. [HSS24b, § 2.3]).
@@
    If $\dim(N\setminus N_{\operatorname{fi}})\leq n-d$, then $N$ is a complete intersection and hence Cohen-Macaulay. In particular $N$ is non-singular at $x\in X$ if and only if $x\in N_{\operatorname{fi}}$.
@@
For any point $x\in X$, the condition (H2) implies that for any $v\in T_{x}X$ and any $\xi\in\mathfrak{g}$, we have
@@
is $(\mathfrak{g}/\mathfrak{g}_{x})^{*}\subset\mathfrak{g}^{*}$. Thus the tangent map $d_{x}\mu$ is surjective if and only if $G_{x}$ is finite. In particular, since $N_{\operatorname{fi}}$ is non-empty, the map $\mu$ is dominant.
For (2.6.1), since $\mu$ is dominant, the shell $N$ is a complete intersection if and only if $N$ has pure dimension ndn-d by [Har77, Theorem 8.21A(c)]. By our assumption and Krull’s Hauptidealsatz, it remains to prove that $\dim(N_{\operatorname{fi}})=n-d$. Note that $X$ is non-singular along $N_{\operatorname{fi}}$ and $\mu$ is smooth along $N_{\operatorname{fi}}$, it follows from [Har77, Proposition 10.4] that $N_{\operatorname{fi}}$ is non-singular of pure dimension ndn-d. So $N$ is Cohen-Macaulay by [Har77, Theorem 8.21A(d)]. Now, it follows from Jacobian criterion that $N$ is non-singular at $x$ only if $X$ is non-singular at $x$ and hence if and only if $x\in N\cap X_{\operatorname{reg}}$ and $\mu$ is smooth at $x$, which is equivalent to that $x\in N_{\operatorname{fi}}$.
For (2.6.2), since the open subset $N_{\operatorname{fi}}$ is dense in $N$ by assumption, the shell $N$ is generically reduced and hence reduced as it is Cohen-Macaulay.
@@ 3.1. Hamiltonian slice theorem
For a scheme $Z$, we denote by $Z^{\operatorname{red}}$ its associated reduced scheme. Let $Y$ be an affine scheme over $\operatorname{Spec}\mathbb{C}$ equipped with a $G$-action. We denote by $Y/\!/G\coloneqq\operatorname{Spec}(\mathcal{O}(Y)^{G})$ the invariant-theoretical quotient with the natural projection $\pi_{Y}\colon Y\rightarrow Y/\!/G$. A subset of $Y$ is $G$-saturated if it is union of fibers of $\pi_{Y}$. We recall the following basic definition:
@@
We require the following Hamiltonian slice theorem of Jung [Jun09], which is the symplectic analogue of Luna’s slice theorem. A simplified proof of this theorem can be found in [BLLT17, § 6] (see also [HSS20, Theorem 3.15] for the linear case). Although the statement in [BLLT17, Theorem 6.1] differs slightly from the one below, the same argument yields our version. For completeness, we include a detailed proof in Appendix § A with some minor adaptations to the details.
@@ 3.2. Isotropy-type stratification
We denote by $(G_{y})$ the conjugacy class of $G_{y}$, $y\in Y$, which we call an isotropy class of $Y$, and an isotropy class is called closed if the orbit $O_{y}=Gy$ is closed in $Y$. Then a closed isotropy class is reductive by Matsushima’s theorem and there are only finitely many closed isotropy classes of $Y$ by Luna’s slice theorem [Lun73]. For a closed isotropy class $(H)$ of $G$, we denote by $(Y/\!/G)_{(H)}$ the points in Y//GY/\!/G corresponding to closed orbits with isotropy class in $(H)$. This is a constructible subset of Y//GY/\!/G by Luna’s slice theorem. Let $Y^{(H)}$ be the inverse image of $(Y/\!/G)_{(H)}$ in $Y$.
Assume in addition that Y//GY/\!/G is irreducible, then there is a unique closed isotropy class $(H)$, called the principal isotropy class, such that $(Y/\!/G)_{\operatorname{pr}}\coloneqq(Y/\!/G)_{(H)}$ is dense in Y//GY/\!/G and the closed orbits $O_{y}$ with $G_{y}\in(H)$ are called principal orbits. Then $(Y/\!/G)_{\operatorname{pr}}\subset Y/\!/G$ is called the principal stratum and it is open in Y//GY/\!/G by Luna’s slice theorem. Let $Y_{\operatorname{pr}}\coloneqq\pi_{Y}^{-1}((Y/\!/G)_{\operatorname{pr}})$. We say that $Y$ has finite principal isotropy groups (FPIG) if its principal isotropy groups are finite.

### v2

@@ Abstract.
Let $Y$ be an irreducible non-singular affine $G$-variety with a $2$-large action. We show that the Hamiltonian reduction $T^{*}Y/\!\!/\!\!/G$ is a symplectic variety with terminal singularities, isomorphic to the affine closure of $T^{*}Z_{\operatorname{reg}}$ where $Z\coloneqq Y/\!/G$. Furthermore, we provide sufficient conditions for the non-existence of a symplectic resolution for such varieties. These results yield three main applications: (i) providing a short proof of G. Schwarz’s theorem on the graded surjectivity of the push-forward map $\mathcal{D}(Y)^{G}\to\mathcal{D}(Z)$; (ii) establishing the surjectivity of the symbol map on $Z$; and (iii) confirming the non-linear analog of a conjecture of Kaledin–Lehn–Sorger for $2$-large actions.
@@ 1. Introduction
Recently, we started a systematic investigation of a geometric construction of symplectic varieties through cotangent bundles. Namely for an irreducible non-singular variety $Y$ with $T^{*}Y$ its cotangent bundle, denote by $\mathcal{O}(T^{*}Y)$ the ring of regular functions on $T^{*}Y$. Then the affine closure $\mathcal{Z}_{Y}$ of $T^{*}Y$ is defined as the normal affine scheme ${\rm Spec}(\mathcal{O}(T^{*}Y))$ so that we have the natural affinization morphism
@@
When $Y$ is a non-singular projective variety, it was shown in [FL25b] that $\mathcal{Z}_{Y}$ is a symplectic variety if the tangent bundle $T_{Y}$ is big and if the projective bundle $\mathbb{P}T_{Y}$ is of Fano type. The most interesting case is where $Y$ is quasi-affine but not affine. In this case, $\psi_{Y}$ is just an open embedding and, therefore, $\mathcal{Z}_{Y}$ is the affine closure of $T^{*}Y$. It remains a challenging problem to determine when $\mathcal{Z}_{Y}$ is a symplectic variety.
Another geometric way to construct symplectic varieties is through Hamiltonian reductions, which is a powerful tool in both physics and geometry to produce new symplectic manifolds from a Hamiltonian action on a symplectic manifold. To be more precise, let $(X,\omega)$ be a non-singular affine symplectic variety on which there exists a Hamiltonian action by a reductive complex linear algebraic group $G$ so that the associated moment map $\mu:X\to\mathfrak{g}^{*}$ is $G$-equivariant (see § 2). We denote by $N_{X}$ the scheme-theoretic fiber $\mu^{-1}(0)$, called the shell. The Hamiltonian reduction X///GX/\!\!/\!\!/G is by definition the affine GIT quotient:
@@
Let $G$ be a reductive group and $Y$ an irreducible non-singular affine $G$-variety that is $2$-large. Then the following holds:
@@
    The Hamiltonian reduction $T^{*}Y/\!\!/\!\!/G$ is the affine closure of $T^{*}(Y/\!/G)_{\rm reg}$. In particular, if Y//GY/\!/G is non-singular, then $T^{*}Y/\!\!/\!\!/G\cong T^{*}(Y/\!/G)$.
@@
    The Hamiltonian reduction $T^{*}Y/\!\!/\!\!/G$ is non-singular if and only if Y//GY/\!/G is non-singular.
The $2$-largeness condition in Theorem 1.1 is not overly restrictive: as shown in [Sch95, Corollary 11.6 (2)] (see also [HSS20, Theorem 3.6]), most $G$-modules are in fact $2$-large, with only finitely many exceptions under mild assumptions. On the other hand, for the special case in which $Y$ is a linear representation of $G$, part (1.1.1) of Theorem 1.1 was proved in [HSS20] (see also [HSS24a, HSS24b] for other related results).
@@
Let $G$ be a reductive group and $Y$ an irreducible non-singular affine $G$-variety that is $2$-large. Set $Z\coloneqq Y/\!/G$ with the quotient map $\pi_{Y}\colon Y\rightarrow Z$.
@@
We next investigate the ($\mathbb{Q}$-)factoriality of the Hamiltonian reduction $T^{*}Y/\!\!/\!\!/G$ and use this to deduce the non‑existence of symplectic resolutions under certain conditions. Recall that the abelianization of $G$ is $G_{\operatorname{ab}}\coloneqq G/[G,G]$. A group $G$ is called perfect if $G_{\operatorname{ab}}=\{e\}$, i.e., $G=[G,G]$.
Theorem 1.3.
Let $G$ be a reductive group and $Y$ an irreducible non-singular affine $G$-variety that is $2$-large. Assume that $\mathcal{O}(Y)^{\times}=\mathbb{C}^{\times}$ or $G$ is connected.
  1. (1.3.1)
    If $G_{\operatorname{ab}}$ is finite and $\operatorname{Pic}(Y)$ is finite, then $\operatorname{Cl}(T^{*}Y/\!\!/\!\!/G)$ is finite. In particular, $T^{*}Y/\!\!/\!\!/G$ is $\mathbb{Q}$-factorial.
  2. (1.3.2)
We present a partial converse of Theorem 1.3 in Corollary 4.8. For $Y$ a $3$-large linear representation of $G$, the local factoriality and $\mathbb{Q}$-factoriality of $T^{*}Y/\!\!/\!\!/G$ were already established in [BS19, Theorem 1.2]. By combining Theorem 1.1 and Theorem 1.3, we obtain the following result.
Theorem 1.4.
Let $G$ be a semisimple group, and let $Y$ be an irreducible non-singular affine $G$-variety which is $2$-large. Assume that $\operatorname{Pic}(Y)$ is finite. Then $T^{*}Y/\!\!/\!\!/G$ admits a symplectic resolution if and only if Y//GY/\!/G is non-singular.
Theorem 1.4 yields a positive answer to a non-linear analog for all $2$-large actions of the following conjecture of D. Kaledin, M. Lehn and C. Sorger, which in the linear case follows from [BS19, Corollary 1.3].
Conjecture 1.5.
Let $G$ be a semisimple group and let $V$ be a finite dimensional representation of $G$. If there exists a symplectic resolution of $T^{*}V/\!\!/\!\!/G$, then V//GV/\!/G is non-singular.
@@ 2. Quasi-Hamiltonian varieties
Throughout this paper, we work over complex numbers, and $G$ always denotes a reductive complex linear algebraic group. By varieties we mean separated reduced schemes of finite type over $\mathbb{C}$, which are not necessarily irreducible.
@@ 2.1. Hamiltonian actions
Then $X$ is called a Hamiltonian $G$-variety so that $V_{\xi}$ and $H_{\xi}$ are called the Hamiltonian vector field and the Hamiltonian function associated with $\xi$, respectively.
For a Hamiltonian action $G$ on $X$, we can define a morphism $\mu_{G,X}\colon X\rightarrow\mathfrak{g}^{*}$ using the following formula:
@@
Given a quasi-Hamiltonian $G$-variety, since $X$ is normal and $\operatorname{codim}(X\setminus X_{\operatorname{reg}})\geq 2$, the moment map $\mu_{G,X_{\operatorname{reg}}}\colon X_{\operatorname{reg}}\rightarrow\mathfrak{g}^{*}$ extends uniquely to a $G$-equivariant morphism $X\rightarrow\mathfrak{g}^{*}$, again denoted by $\mu_{G,X}$. We shall call $\mu_{G,X}$ also the moment map of the quasi-Hamiltonian $G$-variety $X$.
@@
Let $Y$ be an irreducible non-singular $G$-variety, $X\coloneqq T^{*}Y$ the cotangent bundle of $Y$. Then $X$ is a non-singular symplectic variety with the canonical symplectic form $d\eta$, where $\eta$ is the Liouville $1$-form on $T^{*}Y$. Moreover, the lifted action of $G$ on $X$ is Hamiltonian, and the moment map is given by
@@
where $y\in Y$, $\alpha\in T_{y}^{*}Y$ and $\xi\in\mathfrak{g}$. More generally, singular affine Hamiltonian $G$-varieties can also be constructed by considering the affine closure of cotangent bundles. We refer the reader to [FL25a, FL25b] for examples.
@@ 2.2. The shell
Let $X$ be a quasi-Hamiltonian $G$-variety. The shell $N_{X}$ is defined to be the scheme-theoretic fiber of the moment map $\mu\colon X\rightarrow\mathfrak{g}^{*}$ over $0\in\mathfrak{g}^{*}$. We denote by $N_{X,\operatorname{fi}}$ the subset of $N_{X}$ consisting of points $x\in N_{X}\cap X_{\operatorname{reg}}$ such that the isotropy subgroup $G_{x}$ is finite. If $X$ is clear from the context, we will simply denote $N_{X}$ by $N$ and $N_{X,\operatorname{fi}}$ by $N_{\operatorname{fi}}$. The following result is well known for experts, we include a complete proof for the reader’s convenience. See also [KLS06, Lemmas 3.2 and 3.3] and [HSS24b, § 2.3].
Lemma 2.6.
@@
    If $\dim(N\setminus N_{\operatorname{fi}})\leq n-d$, then $N$ is a complete intersection, and hence Cohen-Macaulay. In particular $N$ is non-singular at $x\in X$ if and only if $x\in N_{\operatorname{fi}}$.
@@
For any point $x\in X$, Condition (H2) implies that for any $v\in T_{x}X$ and any $\xi\in\mathfrak{g}$, we have
@@
is $(\mathfrak{g}/\mathfrak{g}_{x})^{*}\subset\mathfrak{g}^{*}$. Thus, the tangent map $d_{x}\mu$ is surjective if and only if $G_{x}$ is finite. In particular, since $N_{\operatorname{fi}}$ is non-empty, the map $\mu$ is dominant.
For (2.6.1), since $\mu$ is dominant, the shell $N$ is a complete intersection if and only if $N$ has pure dimension ndn-d by [Har77, Theorem 8.21A(c)]. By our assumption and Krull’s Hauptidealsatz, it remains to prove that $\dim(N_{\operatorname{fi}})=n-d$. Note that $X$ is non-singular along $N_{\operatorname{fi}}$ and $\mu$ is smooth along $N_{\operatorname{fi}}$, it follows from [Har77, Proposition 10.4] that $N_{\operatorname{fi}}$ is non-singular of pure dimension ndn-d. So $N$ is Cohen-Macaulay by [Har77, Theorem 8.21A(d)]. Now, it follows from the Jacobian criterion that $N$ is non-singular at $x$ only if $X$ is non-singular at $x$ and hence if and only if $x\in N\cap X_{\operatorname{reg}}$ and $\mu$ is smooth at $x$, which is equivalent to $x\in N_{\operatorname{fi}}$.
For (2.6.2), since the open subset $N_{\operatorname{fi}}$ is dense in $N$ by assumption, the shell $N$ is generically reduced and therefore reduced as it is Cohen-Macaulay.
@@ 3.1. Hamiltonian slice theorem
For a scheme $Z$, we denote by $Z^{\operatorname{red}}$ its associated reduced scheme. Let $Y$ be an affine scheme over $\operatorname{Spec}\mathbb{C}$ equipped with a $G$-action. We denote by $Y/\!/G\coloneqq\operatorname{Spec}(\mathcal{O}(Y)^{G})$ the invariant-theoretical quotient with the natural projection $\pi_{Y}\colon Y\rightarrow Y/\!/G$. A subset of $Y$ is $G$-saturated if it is a union of fibers of $\pi_{Y}$. We recall the following basic definition:
@@
We require the following Hamiltonian slice theorem of Jung [Jun09], which is the symplectic analog of Luna’s slice theorem. A simplified proof of this theorem can be found in [BLLT17, § 6] (see also [HSS20, Theorem 3.15] for the linear case). Although the statement in [BLLT17, Theorem 6.1] differs slightly from the one below, the same argument yields our version. For completeness, we include a detailed proof in Appendix § A with some minor adaptations to the details.
@@ 3.2. Isotropy-type stratification
We denote by $(G_{y})$ the conjugacy class of $G_{y}$, $y\in Y$, which we call an isotropy class of $Y$, and an isotropy class is called closed if the orbit $O_{y}=Gy$ is closed in $Y$. Then a closed isotropy class is reductive by Matsushima’s theorem, and there are only finitely many closed isotropy classes of $Y$ by Luna’s slice theorem [Lun73]. For a closed isotropy class $(H)$ of $G$, we denote by $(Y/\!/G)_{(H)}$ the points in Y//GY/\!/G corresponding to closed orbits with isotropy class in $(H)$. This is a constructible subset of Y//GY/\!/G by Luna’s slice theorem. Let $Y^{(H)}$ be the inverse image of $(Y/\!/G)_{(H)}$ in $Y$.
Assume in addition that Y//GY/\!/G is irreducible, then there is a unique closed isotropy class $(H)$, called the principal isotropy class, such that $(Y/\!/G)_{\operatorname{pr}}\coloneqq(Y/\!/G)_{(H)}$ is dense in Y//GY/\!/G and the closed orbits $O_{y}$ with $G_{y}\in(H)$ are called principal orbits. Then $(Y/\!/G)_{\operatorname{pr}}\subset Y/\!/G$ is called the principal stratum and is open in Y//GY/\!/G by Luna’s slice theorem. Let $Y_{\operatorname{pr}}\coloneqq\pi_{Y}^{-1}((Y/\!/G)_{\operatorname{pr}})$. We say that $Y$ has finite principal isotropy groups (FPIG) if its principal isotropy groups are finite.